OpenStudy (anonymous):

a 42-km/h wind blows toward 215degrees, while a plane heads toward 125degrees at 152km/h. what is the resultant velocity of the plane? PLEASE HELP ME

OpenStudy (mayankdevnani):

Use the Law of Cosines to solve this problem, i.e., A^2 = B^2 + C^2 - 2(B)(C)(cos a) where A = resultant velocity of plane B = wind speed = 42 kph (given) C = plane speed = 152 kph (given) a = angle between the plane and wind vectors = 215 - 125 = 90 degrees Since the angle "a" = 90 degrees, then the effective formula simply becomes the Pythagorean theorem, i.e., A^2 = B^2 + C^2 and substituting values, A^2 = 42^2 + 152^2 and solving for A, A = 157.50 kph

OpenStudy (mayankdevnani):

ok @samancha85

OpenStudy (anonymous):

ok but we are doing vectors and magnitudes with this assignment